OpenStudy (anonymous):

A circle has a radius of 7. Find the area of the circle to the nearest whole number.

OpenStudy (raden):

use the formula : area = pi * r^2 with pi = 22/7 and given r = 7

OpenStudy (anonymous):

Formula is- Pi times (radius squared) That gives us, 3.14* 7*7 =3.14*49 =153.86 ( rounded to 154 square units)